# Can LLMs Beat Classical Hyperparameter Optimization Algorithms? A Study on autoresearch

### 25 Mar 2026 v1

@@ A.1 Convergence by Trial Number
The main text reports convergence against cumulative training time, which is our primary comparison. Figures˜3 and 4 show convergence by trial number instead, measuring sample efficiency. These views can differ substantially because LLM-based methods spend additional time on inference between trials, compressing their wall-time curves even when they are competitive per trial.
@@
Figure 4: 0.8B vs 27B by trial number (mean $\pm$ std across 3 seeds). Same methods as Figure˜2. Solid lines: 27B, dashed: 0.8B.
A.2 Frontier Model Comparison: Gemini 2.5 Flash
To test whether a stronger LLM optimizer changes the balance between classical and LLM-based methods, we ran Centaur, Karpathy Agent (Code), and LLAMBO (Paper) with Gemini 2.5 Flash [Comanici et al., 2025] as the LLM optimizer instead of Qwen3.5-27B. Figure˜5 compares the two over cumulative training time. Classical methods (TPE, Random) are shown as references. Within this window, Gemini 2.5 Flash does not outperform Qwen3.5-27B on this task, suggesting that simply swapping to a frontier model does not automatically close the gap to classical methods.
@@
A.3 Incumbent Traces
Beyond aggregate convergence, we visualize per-method incumbent traces to reveal when and how each optimizer finds improvements. Figures˜6 and 7 show incumbent traces against cumulative training time (hours) for all methods. Grey dots are all trials, colored dots are new incumbents, and the staircase line is the best-so-far trajectory. Each panel shows the best seed for that method.
@@ A.3 Incumbent Traces
A.4 Qualitative Agent Behavior
@@ A.5 Diversity Metrics
A.6 Centaur LLM Ratio Ablation
We ablate the fraction of trials delegated to the LLM in Centaur. Table˜4 reports best val_bpb for each ratio on seed 0. The default ratio of r=0.3r{=}0.3 balances classical and LLM contributions; higher ratios degrade performance, particularly for the 27B model where r=0.8r{=}0.8 performs worse than CMA-ES alone. This confirms that CMA-ES needs to retain majority control of the optimization trajectory, with the LLM contributing occasional informed suggestions rather than dominating the search.
Table 4: Centaur LLM ratio ablation (seed 0). Best val_bpb for each ratio. Bold = best per column. The default ratio is r=0.3r{=}0.3.
LLM Ratio $r$ Centaur [0.8B] Centaur [27B]
0.1 0.9746 0.9752
0.2 0.9735 0.9753
0.3 (default) 0.9776 0.9756
0.5 0.9787 0.9745
0.8 0.9789 0.9903
Figure 8: Effect of LLM ratio on Centaur performance (seed 0). Higher ratios give the LLM more trials. CMA-ES alone (dotted) shown as reference. Too much LLM control degrades performance, especially for the 27B model at r=0.8r{=}0.8.

### 29 Mar 2026 v2

@@ A.1 Convergence by Trial Number
The main text reports convergence against cumulative training time, which is our primary comparison. Figures˜4 and 5 show convergence by trial number instead, measuring sample efficiency. These views can differ substantially because LLM-based methods spend additional time on inference between trials, compressing their wall-time curves even when they are competitive per trial.
@@
Figure 5: 0.8B vs 27B by trial number (mean $\pm$ std across 3 seeds). Same methods as Figure˜2.
A.2 Frontier Model Comparison
To test whether a stronger LLM optimizer changes the balance between classical and LLM-based methods, we ran Karpathy Agent (Code) and Centaur with three Gemini variants: Gemini 2.5 Flash [Comanici et al., 2025], Gemini 3.1 Flash-Lite [Google DeepMind, 2026], and Gemini 3.1 Pro Preview [Google DeepMind, 2026]. Gemini 2.5 Flash and 3.1 Flash-Lite runs were stopped early after 16–18 hours of cumulative training time as no meaningful improvement was observed beyond that point. Gemini 3.1 Pro Preview was run for the full 24-hour budget but limited to a single seed due to API cost constraints.
[…]
@@ A.2 Frontier Model Comparison
Figure 7: Gemini 3.1 Flash-Lite vs Qwen3.5-27B as LLM optimizer (cumulative training time). Solid: Gemini 3.1, dashed: Qwen3.5-27B. Same color per method.
A.3 Centaur LLM Ratio Ablation
We ablate the fraction of trials delegated to the LLM in Centaur. Table˜4 reports best val_bpb for each ratio. The default ratio of r=0.3r{=}0.3 balances classical and LLM contributions; higher ratios degrade performance, particularly for the 27B model where r=0.8r{=}0.8 performs worse than CMA-ES alone. This confirms that CMA-ES needs to retain majority control of the optimization trajectory, with the LLM contributing occasional informed suggestions rather than dominating the search.
[…]
LLM Ratio $r$ Centaur [0.8B] Centaur [27B]
[…]
Figure 8: Effect of LLM ratio on Centaur performance (2 seeds, 3 for default). Higher ratios give the LLM more trials. CMA-ES alone (dotted) shown as reference. Too much LLM control degrades performance, especially for the 27B model at r=0.8r{=}0.8.
@@ A.3 Centaur LLM Ratio Ablation
A.4 Incumbent Traces
Beyond aggregate convergence, we visualize per-method incumbent traces to reveal when and how each optimizer finds improvements. Figures˜9 and 10 show incumbent traces against cumulative training time (hours) for all methods. Grey dots are all trials, colored dots are new incumbents, and the staircase line is the best-so-far trajectory. Each panel shows the best seed for that method.
@@ A.4 Incumbent Traces
A.5 Qualitative Agent Behavior